OpenStudy (anonymous):

What would be the equation of $186 in the bank with an interest rate of .001% in year 1, 2, 3, 4, and 5

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@johnweldon1993

jimthompson5910 (jim_thompson5910):

0.001% = 0.001/100 = 0.00001 If you're dealing with simple interest, then y = 0.00001*186x = 0.00186x where x = year number y = amount in the account in year x

jimthompson5910 (jim_thompson5910):

btw 0.001% interest is so very very small. Are you sure that it says that?

OpenStudy (anonymous):

yes. this is the current bank interest for savings. it's super low

OpenStudy (anonymous):

so after the first year, the person would have 186.00186 in their account if they didn't touch it?

jimthompson5910 (jim_thompson5910):

oh sorry I meant y = 186+0.00001*186x = 186+0.00186x

jimthompson5910 (jim_thompson5910):

but yeah, when x = 1, y = 186.00186

OpenStudy (anonymous):

okay, and the second year would be

OpenStudy (anonymous):

like how do u find it?

jimthompson5910 (jim_thompson5910):

by plugging in x = 2

OpenStudy (anonymous):

OH

OpenStudy (anonymous):

so the second year would be 186.0032?

jimthompson5910 (jim_thompson5910):

186+0.00186x = 186+0.00186*2 = 186.00372

jimthompson5910 (jim_thompson5910):

do the same for x = 3 through x = 5

OpenStudy (anonymous):

okay thank you so much!